• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Norum

    Log entry time 19:01:39 on May10,2011

    Entry number 358844

    keyword=Tuesday Day Shift Summary

    Morning: Jack Segal has made temporary repairs to the L-HRS Q1 controller. It is now fixed in negative polarity which is fine for now.
    Beam studies from 9 on
    Vince checked on some timing issues in the hall during beam studies. Separate HALOG re: that.